# The Riemann Hypotheses is false

### v27

@@
 [^1] Data availability statement: My manuscript has no associated data. Introduction. The Riemann zeta function is a function of the complex variable $s$, defined in the half-plane $\Re(s)>1$ by $\zeta(s):=\sum_{n=1}^{\infty}n^{-s}$ and in the whole complex plane by holomorphic continuation. Euler noticed that for $\Re(s)>1$, $\zeta(s)$ can be expressed as a product $\prod_{p}(1-p^{-s})^{-1}$ over the entire set of primes, which entails that $\zeta(s)\neq 0$ for $\Re(s)>1$. It can be shown that $\zeta(s)$ extends to $\mathbb{C}$ as a meromorphic function with only a simple pole at $s=1$, with residue $1$, and satisfies the functional equation $\xi(s)=\xi(1-s)$, where $\xi(z)=\frac{1}{2}z(z-1)\pi^{-z/2}\Gamma(\frac{1}{2}z)\zeta(z)$ and $\Gamma(w)=\int_{0}^{\infty}e^{-x}x^{w-1}\mathrm{d}x$. Let $\chi$ be a primitive Dirichlet character and $L(s,\chi)$ be the corresponding L-function. Most of the analytic properties for $\zeta$ are also known to hold for $L(s,\chi)$ [5, Chapter 10]. Let $\Theta_{\chi}$ be the supremum of the real parts of the zeros of $L(s,\chi)$. The Generalised Riemann Hypothesis (GRH) is equivalent to the statement that $\Theta_{\chi}=\frac{1}{2}$ for each $\chi$, and the Generalised Prime Number Theorem is equivalent to the statement that $L(s,\chi)\neq 0$ at $\Re(s)=1$ for each $\chi$. For a far more thorough discussion of the RH and GRH, the interested reader is kindly referred to Borwein et al [1].
@@
$\sum_{\frac{T}{\log^{2}T}<n<T}n^{\sigma-1+it}\ll T^{\sigma-1+it}\ll T^{\sigma-\frac{1}{2}}\log^{3}T$ (6)
@@
$\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\frac{T^{\frac{1}{2}}\log^{4}T}{m^{1+\frac{1}{\log T}}\log(T/m)}$ (11)
@@
$\sum_{\frac{T}{\log^{2}T}<n<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\frac{T^{\frac{1}{2}}\log^{4}T}{m^{1+\frac{1}{\log T}}\log(T/m)}.$ (12)

### v28

@@
 [^1] Data availability statement: This manuscript has no associated data. Introduction. The Riemann zeta function is a function of the complex variable $s$, defined in the half-plane $\Re(s)>1$ by $\zeta(s):=\sum_{n=1}^{\infty}n^{-s}$ and in the whole complex plane by holomorphic continuation. Euler noticed that for $\Re(s)>1$, $\zeta(s)$ can be expressed as a product $\prod_{p}(1-p^{-s})^{-1}$ over the entire set of primes, which entails that $\zeta(s)\neq 0$ for $\Re(s)>1$. It can be shown that $\zeta(s)$ extends to $\mathbb{C}$ as a meromorphic function with only a simple pole at $s=1$, with residue $1$, and satisfies the functional equation $\xi(s)=\xi(1-s)$, where $\xi(z)=\frac{1}{2}z(z-1)\pi^{-z/2}\Gamma(\frac{1}{2}z)\zeta(z)$ and $\Gamma(w)=\int_{0}^{\infty}e^{-x}x^{w-1}\mathrm{d}x$. Let $\chi$ be a primitive Dirichlet character and $L(s,\chi)$ be the corresponding L-function. Most of the analytic properties for $\zeta$ are also known to hold for $L(s,\chi)$ [5, Chapter 10]. Let $\Theta_{\chi}$ be the supremum of the real parts of the zeros of $L(s,\chi)$. The Generalised Riemann Hypothesis (GRH) is equivalent to the statement that $\Theta_{\chi}=\frac{1}{2}$ for each $\chi$, and the Generalised Prime Number Theorem is equivalent to the statement that $L(s,\chi)\neq 0$ at $\Re(s)=1$ for each $\chi$. For a far more thorough discussion of the RH and GRH, the interested reader is kindly referred to Borwein et al [1].
@@
$\sum_{\frac{T}{\log^{2}T}<n<T}n^{\sigma-1+it}\ll T^{\sigma-\frac{1}{2}}\log^{3}T$ (6)
@@
$\sum_{\frac{T}{\log^{2}T}<n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\frac{T^{\frac{1}{2}}\log^{4}T}{m^{1+\frac{1}{\log T}}|\log(T/m)|}$ (11)
@@
$\sum_{\frac{T}{\log^{2}T}<n<T}\frac{n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}\ll\frac{T^{\frac{1}{2}}\log^{4}T}{m^{1+\frac{1}{\log T}}|\log(T/m)|}.$ (12)
@@
Remark 3.
(i) From the literature on arXiv and elsewhere, it doesn’t appear like this approach on RH has been tried before. However, the idea of applying Gonek’s formula (16) in estimating sums over zeros of zeta in general, is not new. As one can confirm from the literature, it has proven to be a useful tool for such tasks. It is therefore instructive not to rush to bound the summation in (13), e.g. by applying $n^{it}\ll 1$. Doing so will definitely make one lose important information needed for a sharper bound for the sum over $n$.
(ii) The key part of our argument is the estimate in (12), whose derivation is a routine application of partial summation from (5). It actually suffices for the reader to only verify this part. Infact for our purposes, the bound in (12) only has to be $\ll\frac{T}{m^{1+\frac{1}{\log T}}(\log^{3}T)|\log(T/m)|}$.